OpenStudy (anonymous):

Solve the system of equations. Show your work to help me understand, please? x + 2y = 5 3x + 2y = 17

OpenStudy (anonymous):

okay you see 2y in both equations, what can you do?

OpenStudy (anonymous):

subsitution method? I'm honestly not too sure, im bad at this

OpenStudy (anonymous):

Subtract 2y so it will say x=-2y+5 Plug in -2y+ 5 for x in 3x+2y=17 So 3(-2y+5)+2y=17 Rest should be easy

jimthompson5910 (jim_thompson5910):

you can also use the elimination method to subtract equation 2 from equation 1 to get x + 2y = 5 3x + 2y = 17 ------------- -2x + 0y = -12 -2x = -12

OpenStudy (nathan917):

1. Simplifying x + 2y = 5 Solving x + 2y = 5 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2y' to each side of the equation. x + 2y + -2y = 5 + -2y Combine like terms: 2y + -2y = 0 x + 0 = 5 + -2y x = 5 + -2y Simplifying x = 5 + -2y 2. Simplifying 3x + 2y = 17 Solving 3x + 2y = 17 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-2y' to each side of the equation. 3x + 2y + -2y = 17 + -2y Combine like terms: 2y + -2y = 0 3x + 0 = 17 + -2y 3x = 17 + -2y Divide each side by '3'. x = 5.666666667 + -0.6666666667y Simplifying x = 5.666666667 + -0.6666666667y hope this helps!

jimthompson5910 (jim_thompson5910):

now solve -2x = -12 for x once you have the value of x, you can use this to find y

OpenStudy (nathan917):

@mirandabell902 is this a good explanation?

OpenStudy (anonymous):

Yes, thank you

OpenStudy (nathan917):

If you have anymore questions fan me pls.